# Hardware-Aware Parallel Prompt Decoding for Memory-Efficient Acceleration of LLM Inference

### v2

@@ 4.2 Construction Algorithm
For a fixed tree size , we explore all combinations of and where , to identify the dynamic sparse tree that maximizes . To determine the optimal tree size , we define two key functions: 1. Acceptance length (hardware-independent) and 2. Forward pass latency (hardware-dependent). The speedup ratio, , is estimated using a validation dataset, with evaluated once and tested on different hardware types. We aim to find the value of that maximizes .
@@ 5 Experiments
Models and testbeds. We conducted all the experiments using MobileLLaMA-1.4B [6], Vicuna-7B and Vicuna-13B [5]. We used 3 prompt tokens and 1 EPT per prompt token for all inference experiments. The inference throughputs of the models are evaluated on a single NVIDIA A100 GPU with 40GB of memory and a GeForce RTX 4090 using a batch size of 1 and FP16 precision. Further details about the experimental setup can be found in Appendix C.
Training. We froze all trainable parameters of the original LLM. Prompt token embeddings were trained using distillation logits generated from the ShareGPT dataset [22], with a maximum context length of 1024, a cosine learning rate scheduler starting at 0.01, and no warmup. Prompt token embeddings are initialized with normal text token embeddings.
Datasets. We assess the throughput performance of PPD across various tasks and datasets. Specifically, we evaluated PPD using the MT-Bench dataset [25], which contains multi-turn questions with a range of topics, in both non-greedy (temperature follows the default configuration) and greedy settings (temperature=0). We used the GSM8K [7] and HumanEval [3] datasets only in the greedy setting. The GSM8K dataset consists of grade school math problems and we used the first 500 questions of the test split for our evaluations. HumanEval includes coding tasks, for which we set a maximum new token limit of 512 to control the length of the generated sequences. We used the Alpaca [15] dataset as the validation dataset to produce the latencies and acceptance lengths used for dynamic sparse tree construction.
@@ 5.1 Speedup Comparison with Parallel Decoding Methods
Figure 5 displays the throughput of PPD on MT-Bench, HumanEval, and GSM8K with temperature equal to 0. PPD achieves consistent walltime speedup ratios from 2.12 to 2.49 on different GPUs. In general, PPD performs better in coding and math reasoning tasks, achieving speedups between 2.21 and 2.49. This can be attributed to the fact that both code and math equations often contain fixed patterns and repetitive symbols, which narrows the range of plausible candidates and simplifies the prediction. We also found that with typical acceptance, the speedup increases with temperature. Another notable trend is that smaller models, such as Vicuna-7B, generally achieve more significant speedup ratios as compared to larger models, like Vicuna-13B. PPD aims to generate more tokens per step, which comes with increased computational demands. For larger models that already require substantial computational resources, it is necessary to limit the size of the sparse tree to avoid exceeding the GPU’s utilization cap and causing increased latency. As a result, the number of tokens accepted per step is reduced, leading to lower speedups. However, this can be amortized when using more powerful GPUs than the NVIDIA A100 and the RTX 4090, such as NVIDIA H100.
@@ 5.2 Long-range Token Prediction
5.3 Memory Efficiency and Synergistic Integrations with Speculative Decoding
Memory efficiency. As shown in Figure 7, we compare the memory overhead of PPD with the leading parallel decoding (Medusa) and speculative decoding approaches (Eagle). The memory overhead of PPD is just 0.004% of Medusa’s and 0.007% of Eagle’s. This efficiency stems from the efficient use of embeddings in PPD, which are significantly smaller than decoding heads and draft models, both of which scale with vocabulary size.
PPD + Speculative Decoding. As an orthogonal optimization in accelerating LLMs, PPD can be easily integrated with speculative decoding [11]. To demonstrate this, we applied PPD to Vicuna-68M [24] and used it as the draft model for Vicuna-7B. This combination resulted in a speedup of up to 1.22 for speculative decoding on Vicuna-7B compared to using speculative decoding alone.
@@ 5.4 Ablation Study
Dynamic Sparse Tree. Figure 8(a) shows that dynamic sparse trees consistently achieve longer acceptance lengths compared to static and random ones across varying sizes. The acceptance length for dynamic sparse trees shows a steady increase as the tree size extends, suggesting its good scalability. The convergence of dynamic and static sparse trees at larger sizes suggests a structural similarity emerging from constraints in tree depth and tree node count.
@@ 6 Conclusion
We introduced PPD, a memory-efficient, cost-effective, and powerful parallel decoding method that incorporates a hardware-aware dynamic sparse tree. Utilizing specially trained prompt tokens to predict long-range tokens accurately, PPD achieves a speedup of up to 2.49 in inference while employing only 0.0002% additional trainable parameters and without incorporating new models or architectural components. We believe that PPD offers a novel perspective on the capabilities of parallel decoding. In future work, it could be synergized with other speculative or parallel decoding techniques to expedite inference even further.
Supplementary Material
Hardware-Aware Parallel Prompt Decoding for
Memory-Efficient Acceleration of LLM Inference
Appendix A Training Loss
@@
Appendix B Extended Ablation Study
B.1 Effect of EPTs on Prediction Accuracy
@@
B.2 Impact of Knowledge Distillation (KD), Epochs, and Batch Size on Prediction Accuracy
@@ B.2.2 Knowledge Distillation
B.2.3 Effect of Batch Size
@@
B.3 Prefix Tuning + Prompt Token
Prefix tuning [14], similar to prompt tuning, provides a parameter-efficient approach to fine-tune a pre-trained model. Unlike prompt tuning, it modifies the KV cache of every attention layer by prepending trained vectors. We hypothesize that the combination of prefix tuning and prompt tokens can lead to greater learning capacity and higher prediction accuracy. This hypothesis is based on the intuition that prompt tokens should see a different context than the input tokens when predicting long-range tokens. For example, if the input sequence is "Once upon a time", then enhancing the input with a prompt template might provide more suitable semantic context for long-range prediction. An enhanced input like "Predict the next-next token. Once upon a time" might empower the prompt token to predict the correct next-next token. Prefix tuning serves as the prompt template to enhance the hidden states visible to the prompt tokens.
@@ B.3 Prefix Tuning + Prompt Token
B.4 Custom Decoding Heads + Prompt Token
It has been demonstrated that a fine-tuned decoding head alone can effectively predict long-range tokens [23, 1]. Thus, we hypothesize that combining a separately fine-tuned decoding head with prompt tokens might further enhance the potential of PPD. As shown in Figure 11, we trained a separate decoding head to transform only the hidden states of prompt tokens into logits. A key distinction from Medusa is that this decoding head is responsible for generating tokens at multiple positions, rather than just one.
@@ B.4 Custom Decoding Heads + Prompt Token
B.5 Attention Masking for EPTs
@@
B.5.1 Ensemble Attention Masking
@@
B.5.2 Decoder-like Attention Masking
@@
B.5.3 Encoder-like Attention Masking
@@ B.5.4 Results
B.6 Aggregation Method for EPTs
@@
B.7 Multi-exit Ensemble
@@
Appendix C Experiment Details
@@
In Fig. 2, the temperature settings for PPD, Eagle [16], and Medusa [1] follow the default configuration, while the other models use a greedy setting (temperature=0). This choice is based on findings that retrieval-based methods perform significantly worse in non-greedy settings. Similarly, LOOKAHEAD DECODING [8], REST [9], and PLD [21] in Fig. 4 also use a temperature setting of 0 for the same reasons.
Appendix D Limitations
1.
Low prediction accuracy for very small models. We found that for very small models like Vicuna-68M [24], which only has 2 decoder layers and an embedding dimension of less than 1000, PPD suffers from low prediction accuracy. This is because the embedding dimension determines the expressive power of a prompt token, and the transformer architecture’s depth is crucial for efficient information flow to the prompt tokens.
2.
3.
Appendix E Societal Impact

### v3

@@ 4.2 Two-Stage Tree Pruning Algorithm
To achieve this, we define two key functions: 1. Acceptance length (hardware-independent) and 2. Forward pass latency (hardware-dependent). The speedup ratio, , is estimated using a validation dataset, with evaluated once and tested on different hardware platforms. We then choose the tree size budget that maximizes based on the measured runtime latency on the specific hardware platform. To eliminate runtime overhead, hardware latency profiling is conducted during idle periods.
@@ 5 Experiments
Models and testbeds. We conducted all the experiments using MobileLLaMA-1.4B (Chu et al., 2023), Vicuna-7B and Vicuna-13B (Chiang et al., 2023). We used 3 prompt tokens and 1 EPT per prompt token for all inference experiments. The inference throughputs of the models are evaluated on a single NVIDIA A100 GPU with 40GB of memory and a GeForce RTX 4090 using a batch size of 1 and FP16 precision. Further details about the experimental setup can be found in Appendix H.
Training. We froze all trainable parameters of the original LLM. Prompt token embeddings were trained using distillation logits generated from the ShareGPT dataset (ShareGPT, 2023), with a maximum context length of 1024, a cosine learning rate scheduler starting at 0.01, and no warmup. Prompt token embeddings are initialized with normal text token embeddings. For each model, the same set of prompt tokens is used across experiments to demonstrate its generalizability.
Datasets. We assess the throughput performance of PPD across various tasks and datasets. Specifically, we evaluated PPD using the MT-Bench dataset (Zheng et al., 2023), which contains multi-turn questions with a range of topics, in both non-greedy (temperature follows the default configuration) and greedy settings (temperature=0). We used the GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) datasets only in the greedy setting. The GSM8K dataset consists of grade school math problems and we used the first 500 questions of the test split for our evaluations. HumanEval includes coding tasks, for which we set a maximum new token limit of 512 to control the length of the generated sequences. We used the Alpaca (Li et al., 2023) dataset as the validation dataset to produce the latencies and acceptance lengths used for sparse tree pruning.
@@ 5.1 Speedup Comparison with Parallel Decoding Methods
Figure 5 displays the throughput of PPD on MT-Bench, HumanEval, and GSM8K with temperature equal to 0. PPD achieves consistent walltime speedup ratios from 2.12 to 2.49 on different GPUs, which demonstrates that prompt tokens generalize well on different tasks. In general, PPD performs better in coding and math reasoning tasks, achieving speedups between 2.21 and 2.49. This can be attributed to the fact that both code and math equations often contain fixed patterns and repetitive symbols, which narrows the range of plausible candidates and simplifies the prediction. We also found that with typical acceptance, the speedup increases with temperature. Another notable trend is that smaller models, such as Vicuna-7B, generally achieve more significant speedup ratios as compared to larger models, like Vicuna-13B. PPD aims to generate more tokens per step, which comes with increased computational demands. For larger models that already require substantial computational resources, it is necessary to limit the size of the sparse tree to avoid exceeding the GPU’s utilization cap and causing increased latency. As a result, the number of tokens accepted per step is reduced, leading to lower speedups. However, this can be amortized when using more powerful GPUs than the NVIDIA A100 and the RTX 4090, such as NVIDIA H100.
@@ 5.2 Long-range Token Prediction
5.3 Memory and Training Efficiency
Memory efficiency. As shown in Figure 7(a), we compare the memory overhead of PPD with the leading parallel decoding (Medusa) and speculative decoding approaches (Eagle). The memory overhead of PPD is negligible, whereas Eagle incurs up to 5.8 overhead and Medusa up to 1.6. This efficiency stems from: (1) PPD’s parameter efficiency—token embeddings are much smaller than decoding heads and draft models, which scale with vocabulary size; and (2) using a single KV cache, unlike methods like Eagle that require separate caches for the draft and target models.
| Method | Training Time |
| PPD (Ours) | 0.52 hours |
| Medusa | 1.24 hours |
| Eagle | 1-2 days |
Training efficiency. Table 2 compares the training times of PPD with parallel and speculative decoding methods. PPD is trained until its evaluation accuracy of top-10 candidates surpasses that of Medusa on Alpaca Eval. Notably, PPD surpasses Medusa in evaluation accuracy while training in less than half the time, demonstrating its great potential to reduce training cost.
@@ 5.4 Ablation Study
Tree Attention. As illustrated in Figure 7(b), tree attention boosts the speedup ratio of PPD by an additional 32%, indicating that PPD generates accurate top-k predictions. Even without the use of tree attention, PPD still outperforms all other parallel decoding methods, achieving up to a 14% higher speedup ratio, demonstrating the effectiveness.
Sparse Tree Pruning Algorithm. Figure 8(a) shows that the pruned sparse trees consistently achieve longer acceptance lengths compared to naive and random ones across varying sizes. The acceptance length for pruned sparse trees shows a steady increase as the tree size extends, suggesting its good scalability. The convergence of pruned and naive sparse trees at larger sizes suggests a structural similarity emerging from constraints in tree depth and tree node count.
@@
PPD + Speculative Decoding. As an orthogonal optimization in accelerating LLMs, PPD can be easily integrated with speculative decoding (Kim et al., 2024). To demonstrate this, we applied PPD to Vicuna-68M (Yang et al., 2024) and used it as the draft model for Vicuna-7B. This combination resulted in a speedup of up to 1.22 for speculative decoding on Vicuna-7B compared to using speculative decoding alone.
@@ 6 Conclusion
We introduced PPD, a memory-efficient and cost-effective parallel decoding method that incorporates a hardware-aware online tree optimization. Utilizing specially trained prompt tokens to predict long-range tokens accurately, PPD achieves a speedup of up to 2.49 in inference while employing negligible additional trainable parameters and memory overhead. We showcased that PPD offers a novel perspective on the capabilities of parallel decoding. Importantly, it could be synergized with other speculative or parallel decoding techniques to expedite inference even further.
Limitations
1.
2.
Appendix A Appendix
Appendix B Detailed Tree Construction Algorithm
We follow the same optimal sparse tree construction approach as in Medusa (Cai et al., 2024).
Definition B.1.
Let be the maximal number of prompt tokens per tree node. The sparse tree can exist in states, each represented by corresponding to state , where . Let denote the subtree of composed solely of candidate tokens. The maximum depth of is .
Proposition B.1.
For a sparse tree state , where each candidate token follows a path from the root, and the acceptance probability at each path position , the expected number of tokens generated is given by , where represents the contribution of a token to the expected number of tokens.
Proposition B.2.
The expected total number of tokens generated for the sparse tree state at the current and the next decoding step is given by , where represents the state transition probability from state to state .
We are now ready to introduce Proposition B.3, which we use in the pruning algorithm.
Proposition B.3.
For a sparse tree state with candidate subtree , the change in expected total tokens due to the removal of a prompt token at candidate token is given by , where is the acceptance probability of candidate , denotes the number of prompt tokens prior to removal. We assume that .
Proposition B.4.
The amortized number of tokens generated for the sparse tree state is given by , where is the steady-state probability of state , and is the function defined in Proposition B.1.
The sparse tree construction algorithm can now be formulated as finding the sparse tree with candidate tokens and prompt tokens to maximize :
For a fixed tree size , we explore all combinations of and where , to identify the sparse tree that maximizes .
Appendix C Comparison with More Baselines
C.1 Comparison with Speculative Decoding
| Temperature | Method | Speedup Ratio |
| T=1 (Stochastic) | PPD | 2.26 |
| Eagle | 2.15 | |
| Sps | 1.38 | |
| T=0 (Greedy) | PPD | 2.15 |
| Eagle | 2.60 | |
| Sps | 1.68 |
Table 3 compares the speedup ratios of different speculative decoding methods on Vicuna-7B and MT-Bench. PPD outperforms Eagle at , whereas Eagle achieves better performance at . This trend aligns with findings in typical acceptance Cai et al. (2024), which tends to be more effective at higher temperatures.
PPD serves as a cost-effective and memory-efficient alternative to speculative decoding Li et al. (2024b); Kim et al. (2024), specifically designed for edge and mobile deployments. Both memory usage and training time are critical metrics in this context. More free memory enables the use of larger batch sizes, resulting in improved throughput.
Memory efficiency is also vital for long-context tasks, where a large KV cache size could otherwise lead to OOM errors. Additionally, reducing training time significantly lowers training costs, making PPD more accessible for local deployments. Faster training is particularly valuable in online learning scenarios, where the target model distribution may shift over time, requiring frequent retraining.
C.2 Comparison with More Baselines
| Temperature | Method | Speedup Ratio |
| T = 0.1 | PPD | 2.28 |
| PaSS | 1.47 | |
| T = 0.8 | PPD | 2.40 |
| PaSS | 1.24 |
| Method | Speedup Ratio |
| PPD | 2.26 |
| Draft&Verify | 1.456 |
Appendix D Training Loss
@@
Appendix E Generalizability of Prompt Tokens to Different Tasks
While original prompt tuning tailors LLMs for specific downstream tasks, our prompt tokens are task-agnostic. To demonstrate their generalizability, Figure 10 shows the prediction accuracy of a single set of prompt tokens across three different datasets. Trained on the ShareGPT dataset, these tokens generalize effectively to unseen tasks. We also report the prediction accuracy for using 5 prompt tokens.
Appendix F Extended Ablation Study
F.1 Effect of EPTs on Prediction Accuracy
@@
F.2 Impact of Knowledge Distillation (KD), Epochs, and Batch Size on Prediction Accuracy
@@ F.2.2 Knowledge Distillation
F.2.3 Effect of Batch Size
@@
F.3 Prefix Tuning + Prompt Token
Prefix tuning (Li and Liang, 2021), similar to prompt tuning, provides a parameter-efficient approach to fine-tune a pre-trained model. Unlike prompt tuning, it modifies the KV cache of every attention layer by prepending trained vectors. We hypothesize that the combination of prefix tuning and prompt tokens can lead to greater learning capacity and higher prediction accuracy. This hypothesis is based on the intuition that prompt tokens should see a different context than the input tokens when predicting long-range tokens. For example, if the input sequence is "Once upon a time", then enhancing the input with a prompt template might provide more suitable semantic context for long-range prediction. An enhanced input like "Predict the next-next token. Once upon a time" might empower the prompt token to predict the correct next-next token. Prefix tuning serves as the prompt template to enhance the hidden states visible to the prompt tokens.
@@ F.3 Prefix Tuning + Prompt Token
F.4 Custom Decoding Heads + Prompt Token
It has been demonstrated that a fine-tuned decoding head alone can effectively predict long-range tokens (Stern et al., 2018; Cai et al., 2024). Thus, we hypothesize that combining a separately fine-tuned decoding head with prompt tokens might further enhance the potential of PPD. As shown in Figure 12, we trained a separate decoding head to transform only the hidden states of prompt tokens into logits. A key distinction from Medusa is that this decoding head is responsible for generating tokens at multiple positions, rather than just one.
@@ F.4 Custom Decoding Heads + Prompt Token
F.5 Attention Masking for EPTs
@@
F.5.1 Ensemble Attention Masking
@@
F.5.2 Decoder-like Attention Masking
@@
F.5.3 Encoder-like Attention Masking
@@ F.5.4 Results
F.6 Aggregation Method for EPTs
@@
F.7 Multi-exit Ensemble
@@
Appendix G Effect of Batch Size on Speedup
| Batch Size | 1 | 2 | 3 | 4 |
| PPD Speedup Ratio (w/o Tree Attention) | 1.71 | 1.65 | 1.63 | 1.64 |
| PPD Speedup Ratio (with Tree Attention) | 2.26 | 1.90 | 1.58 | 1.52 |
As shown in Table 13, consistent speedup ratios are achieved across different batch sizes without tree attention. However, with tree attention, the speedup ratio decreases as batch size increases, a pattern similar to other parallel and speculative decoding methods.
Appendix H Experiment Details
@@
In Fig. 2, the temperature settings for PPD, Eagle (Li et al., 2024b), and Medusa (Cai et al., 2024) follow the default configuration, while the other models use a greedy setting (temperature=0). This choice is based on findings that retrieval-based methods perform significantly worse in non-greedy settings. Similarly, LOOKAHEAD DECODING (Fu et al., 2024), REST (He et al., 2023), and PLD (Saxena, 2023) in Fig. 4 also use a temperature setting of 0 for the same reasons.
Appendix I Societal Impact